x/(x-1) + 3/ (x+1) = 7
Answer by stanbon(75887)

x/(x-1) + 3/ (x+1) = 7
Multiply thru by (x-1)(x+1) to get:
x(x+1) + 3(x-1) = 7(x^2-1)
x^2 + x + 3x - 3 = 7x^2 - 7
6x^2 - 4x - 4 = 0
3x^2 - 2x -2 = 0
x = [2 +- sqrt(4 - 4*3*-2)]/6
x = [2 +- sqrt(28)]/6
x = [2 +- 2sqrt(7)]/6
x = (1/3) + (1/3)sqrt(7) or x = (1/3) - (1/3)sqrt(7)
============================================
Cheers,
Stan H.